OpenStudy (anonymous):

Factor completely: 64x2 – 121

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\] in your example \(a=8x, b-11\)

OpenStudy (anonymous):

so which one of these is it Factor completely: 64x2 – 121 Answer (64x – 121)(x + 1) (64x – 121)(x – 1) (8x – 11)(x –11) (8x – 11)(8x + 11

OpenStudy (uri):

(8x-11)(8x+11)

OpenStudy (anonymous):

what about this one Factor completely: 4x2 – 121 Answer (2x – 11)(2x + 11) (2x – 11)(2x – 11) (4x – 121)(x + 1 ) (4x – 121)(x – 1)

OpenStudy (uri):

One question at a time,Close this thread ask in a new one.